# Prot2Token: A Unified Framework for Protein Modeling via Next-Token Prediction

### v1

@@ A.2.3 Sequence-to-sequence
This tokenization is applied when the output is a sequence of labels corresponding residue-by-residue to the input protein sequence, meaning the output token sequence length mirrors the input protein length. Examples include Secondary Structure (SS) prediction, where each amino acid is classified into states like $\alpha$-helix (H), $\beta$-strand (E), or coil (C), forming a target sequence like {H, H, C, …}. Another application is 3D structure prediction using structural alphabets. For instance, the encoder part of a pre-trained VQVAE) model ([39]) converts 3D coordinates into a sequence of discrete 3D_number tokens (e.g., 4096), where each amino acid corresponds to one 3D_i token encoding 3D structural information.
@@ A.3 Dataset
Mutation stability [30] $\approx$1.92 $\dagger$ $\approx$480,000 (5-fold) - Regression
@@ A.3.1 PEER benchmark
The PEER benchmark [27] provides a unified evaluation suite for protein sequence understanding, integrating datasets for protein function, subcellular localisation, secondary structure, protein–protein interaction (PPI), and protein–ligand affinity prediction. Each task is delivered with homology-aware train/validation/test splits and pre-defined evaluation metrics, enabling direct comparison between conventional feature-engineering pipelines, sequence-embedding models, and large-scale protein language models. From PEER we adopt five datasets that align with our experimental focus: (i) human PPI pairs for binary interaction prediction, (ii) secondary-structure assignments for residue-level sequence-to-sequence labelling, (iii) fluorescence intensities for single-sequence regression, (iv) stability ($\Delta$$T_{m}$) measurements for mutation-effect regression, and (v) protein–ligand affinity (PLA) scores for sequence–SMILES binding prediction.
@@ A.3.6 Protein melting temperature
We leveraged the HotProtein [46] sequence-only benchmark to predict protein melting temperatures from primary sequence alone. HotProtein comprises 182,000 amino acid sequences of 230 organisms, each labeled with the optimal growth temperature of its source organism (-20°C to 120°C) as a lower bound proxy of the true melting temperature of the protein. For evaluation, the ProCeSa [51] paper defined 10-fold cross-validation splits on various subsets of the dataset, such as HP-S2C2 (binary: hot vs. cold), HP-S2C5 (five-class), and HP-S (full dataset). In our study, we used only the first fold of the provided splits, further dividing the training portion into training and validation sets.
@@ A.3.16 Protein-protein structure similarity
8 $\text{MG}^{2+}$ Magnesium Ion MG.fasta 2951 9494
@@ A.4.1 Classification
Regarding the Human PPI task, we maintained the ESM model weights as fixed and only unlocked the last four layers of it to be fine-tuned and connected to the decoder. Note that to give the encoder two sequences at one feed for PPI, we concatenated two sequences using the <EOS> token. We observed that adding more tasks helped boost the performance of Human PPI (Table 16). However, Prot2Token tended to overfit on this task, indicating that the improvement from adding auxiliary tasks may be due to the regularization effect of multi-task learning.
@@
We performed dimensionality reduction using t-SNE and UMAP to visualize these embeddings in two dimensions based on their known group assignments. Visualizations revealed that the original pretrained model exhibited weak separation among groups. However, when we repeated the same process using the fine-tuned ESM2 checkpoint (updated only via substrate-based kinase group classification), the resulting projections displayed improved clustering by group. These qualitative trends were confirmed with unsupervised clustering metrics, including the silhouette score and the Calinski-Harabasz index. As shown in Figures 7 and 8 and Table 20, the fine-tuned encoder demonstrates clearer group structure despite never directly observing kinase sequences during training—suggesting that supervised signals from substrates alone can reorganize the encoder’s representation space in a biologically meaningful way.
@@ A.4.2 Regression
In the regression experiments, we fixed the majority of the ESM encoder parameters, unfreezing only the last six layers for joint fine-tuning with the decoder. For comparison, baseline models attach a single linear regression head to a frozen encoder. Because Spearman correlation is invariant to monotonic transformations, we found that min–max scaling the labels to $[0,1]$ with four digits precision after floating point markedly improved convergence and performance: the decoder learns the structure of numerical outputs more rapidly when they occupy a consistent range.
@@
Table 22: Protein–protein binding affinity prediction on PPB-Affinity.
@@
Table 23: Performance of predicting thermostability on HotProtein (HP-S test split, fold 1).
@@ A.4.3 Self-supervised pre-training of decoder
Furthermore, Figures 10 and 11 illustrate the frequency distribution of labels in phosphorylation and protein-ligand binding site tasks, respectively. These figures clearly show the imbalanced and sparse nature of labels, underscoring why explicit inductive biases provided through self-supervised pre-training are crucial for effective model training.
@@ A.4.4 Binding site
Based on the order of ligands presented in Table 24, we grouped the ligands into distinguishable sets of 10, 20, 30, and all 41 ligands. Each ligand in a set was treated as a separate task defined by a task token figure 12, and trained together in one training session. We selected each of those sets and jointly trained them alongside 20 self-supervised tasks using the latest checkpoint from the self-supervised pre-training phase. For this fine-tuning phase, the self-supervised tasks were reduced to a total number of 20k samples. Also, we removed protein samples with lengths greater than 1280 and set batch size to 98,304 tokens. During all training processes, only the last eight blocks of the encoder (ESM2-650m) were fine-tuned, while all non-encoder parameters of the supermodel were fully fine-tuned.
@@
Finally, preliminary experiments were conducted on Protein-Protein Binding Site Prediction using the PPIRef dataset [47]. The task involved predicting the binding interface residues on a target protein given the sequences of both the target and binder proteins (concatenated as input). Binding residues were defined as those on the target protein within 6Å of the binder. Initial results using the simplified index tokenization yielded a positive F1 score of 0.47 on the test set. While promising, this result is preliminary, and further investigation is required.
@@
Ligand Metrics Prot2Token-C TargetS [61] LMetalSite [62] ZinCap [63] MIB2 [59]
@@ A.4.5 Sequence-to-sequence
S-ESM (Prot2Token-C encoder) 44.40 19.34 11.50 0.5806 0.7963
@@
Figure 15: Validation perplexity curve for sequence-to-3D structure prediction. The continued decrease in perplexity throughout training indicates that the model has not yet reached convergence, and additional training may yield further gains in predictive accuracy.
@@ A.5.3 Optimizing feature selection
[…]
@@
[…]
@@
[…]

### v2

@@ A.2.2 Regression
We investigated the impact of token ordering on regression performance by reversing the target sequence to a right-to-left format (least significant digit first). We observed a measurable degradation in performance compared to the standard left-to-right approach. We attribute this to the loss of the coarse-to-fine inductive bias inherent in left-to-right generation, where the model first predicts the most significant digits (establishing magnitude) before refining the value with lower-order precision; reversing this order forces the model to predict fine-grained details without an established context for the overall scalar value, leading to overfitting and reduced accuracy.
@@ A.2.3 Sequence-to-sequence
This tokenization is applied when the output is a sequence of labels corresponding residue-by-residue to the input protein sequence, meaning the output token sequence length mirrors the input protein length. Examples include Secondary Structure (SS) prediction, where each amino acid is classified into states like $\alpha$-helix (H), $\beta$-strand (E), or coil (C), forming a target sequence like {H, H, C, …}. Another application is 3D structure prediction using structural alphabets. For instance, the encoder part of a pre-trained VQVAE) model [39] converts 3D coordinates into a sequence of discrete 3D_number tokens (e.g., 4096), where each amino acid corresponds to one 3D_i token encoding 3D structural information.
@@ A.3 Dataset
[…]
@@ A.3.1 PEER benchmark
[…]
@@ A.3.6 Protein melting temperature
We leveraged the HotProtein [46] sequence‑only benchmark to predict protein melting temperatures from primary sequence alone. HotProtein comprises 182,000 amino acid sequences of 230 organisms, each labeled with the optimal growth temperature of its source organism (-20°C to 120°C) as a lower bound proxy of the true melting temperature of the protein. For evaluation, the ProCeSa [51] paper defined 10-fold cross-validation splits on various subsets of the dataset, such as HP-S2C2 (binary: hot vs. cold), HP-S2C5 (five-class), and HP-S (full dataset). In our study, we used only the first fold of the provided splits, further dividing the training portion into training and validation sets.
@@ A.3.16 Protein-protein structure similarity
8 $\text{MG}^{2+}$ Magnesium Ion MG.fasta 2951 $94$
@@ A.4.1 Classification
Regarding the Human PPI task, we maintained the ESM model weights as fixed and only unlocked the last four layers of it to be fine-tuned and connected to the decoder. Note that to give the encoder two sequences at one feed for PPI, we concatenated two sequences using the <EOS> token. We observed that adding more tasks helped boost the performance of Human PPI (Table 16). However, Prot2Token tended to overfit on this task, indicating that the improvement from adding auxiliary tasks may be due to the regularization effect of multi-task learning. We used early stopping to avoid overfitting.
@@
We performed dimensionality reduction using t-SNE and UMAP to visualize these embeddings in two dimensions based on their known group assignments. Visualizations revealed that the original pretrained model exhibited weak separation among groups. However, when we repeated the same process using the fine-tuned ESM2 checkpoint (updated only via substrate-based kinase group classification), the resulting projections displayed improved clustering by group. These qualitative trends were confirmed with unsupervised clustering metrics, including the silhouette score and the Calinski-Harabasz index. As shown in Figures 6 and 7 and Table 20, the fine-tuned encoder demonstrates clearer group structure despite never directly observing kinase sequences during training—suggesting that supervised signals from substrates alone can reorganize the encoder’s representation space in a biologically meaningful way.
@@ A.4.2 Regression
In the regression experiments, we fixed the majority of the ESM encoder parameters, unfreezing only the last six layers for joint fine-tuning with the decoder. For comparison, baseline models attach a single linear regression head to a frozen encoder. Because Spearman correlation is invariant to monotonic transformations, we found that min–max scaling the labels to $[0,1]$ with four digits precision after floating point markedly improved convergence and performance: the decoder learns the structure of numerical outputs more rapidly when they occupy a consistent range. For the RMSE report, we scaled the numbers back to their original range.
@@
Table 22: Protein–protein binding affinity prediction on PPB-Affinity. The baseline is based on ESM-2 650m encoder.
@@
Table 23: Performance of predicting thermostability on HotProtein (HP-S test split, fold 1). $\dagger$ denotes a linear layer fine-tuned on the last four encoder blocks.
@@ A.4.3 Self-supervised pre-training of decoder
Furthermore, Figures 9 and 10 illustrate the frequency distribution of labels in phosphorylation and protein-ligand binding site tasks, respectively. These figures clearly show the imbalanced and sparse nature of labels, underscoring why explicit inductive biases provided through self-supervised pre-training are crucial for effective model training.
@@ A.4.4 Binding site
Based on the order of ligands presented in Table 24, we grouped the ligands into distinguishable sets of 10, 20, 30, and all 41 ligands. Each ligand in a set was treated as a separate task defined by a task token figure 11, and trained together in one training session.
We selected a discrete task-token representation for ligands, rather than a continuous chemical encoder input (e.g., via SMILES), based on three key considerations. First, empirical evaluations demonstrated that our token-based approach yielded superior performance compared to architectures utilizing current pre-trained chemical encoders, suggesting limitations in the efficacy of chemical encoders for binding residue prediction. Second, this discrete formulation facilitates deep interpretability; by isolating ligand representations, we were able to analyze the learned embedding space to confirm that the model captures meaningful physicochemical relationships between ligands (A.5). Finally, a fixed vocabulary remains highly practical for many applications, such as ion binding, where the target space is naturally bounded, allowing our framework to robustly support a wider range of simultaneous targets (41 types) than many existing specialized predictors.
We selected each of those sets and jointly trained them alongside 20 self-supervised tasks using the latest checkpoint from the self-supervised pre-training phase. For this fine-tuning phase, the self-supervised tasks were reduced to a total number of 20k samples. Also, we removed protein samples with lengths greater than 1280 and set batch size to 98,304 tokens. During all training processes, only the last eight blocks of the encoder (ESM2-650m) were fine-tuned, while all non-encoder parameters of the supermodel were fully fine-tuned.
@@
Finally, preliminary experiments were conducted on Protein-Protein Binding Site Prediction using the PPIRef dataset [47]. The task involved predicting the binding interface residues on a target protein given the sequences of both the target and binder proteins (concatenated as input). Binding residues were defined as those on the target protein within 6Å of the binder. Initial results using the simplified index tokenization yielded an F1 score of 0.47 on the test set. While encouraging, this result is preliminary, and further investigation is required.
@@
Ligand Metrics Prot2Token-C
TargetS
[61]
LMetalSite
[62]
ZinCap
[63]
MIB2
[59]
@@ A.4.5 Sequence-to-sequence
S-ESM (Prot2Token‑C encoder) 44.40 19.34 11.50 0.5806 0.7963
@@
Figure 15: Validation perplexity curve for sequence-to-3D structure prediction. While perplexity keeps decreasing, the validation TM-score saturates at $\approx$0.55 on CAMEO 2024. Post-hoc analysis shows the 3D tokenizer (VQ-VAE) itself reconstructs with an upper bound of $\approx$0.60 TM-score on the same benchmark; hence the plateau reflects a tokenizer-imposed ceiling rather than insufficient optimization of Prot2Token. Improving the tokenizer is likely required to push beyond this regime.
@@ A.5.3 Optimizing feature selection
3:(a) Filter Ligands:    $\text{high\_quality\_ligands}\leftarrow\{\text{ligand}\mid F1(\text{ligand})>\text{threshold}\}$
4:(b) Reduce Embeddings:    $\text{pca\_embeddings}\leftarrow\text{PCA}(\text{task\_token\_embeddings},\,n\,\text{components},\,99\%\ \text{variance})$
@@
6:      $k_{\mathrm{optimal}}\leftarrow\text{ElbowMethod}(\text{pca\_embeddings})$
7:      $\text{target\_clusters}\leftarrow\text{KMeans}(\text{pca\_embeddings},\,k_{\mathrm{optimal}})$ Step 2. Feature Combination Evaluation:
8:      $\text{best\_ari}\leftarrow$-$\infty$
@@
10:  $\text{combinations}\leftarrow\text{Combinations}(\text{features\_pool},\,n_{\mathrm{features}})$
11:  for each $\text{combination}\in\text{combinations}$ do
12:    $\text{feature\_clusters}\leftarrow\text{KMeans}(\text{combinations},\,k_{\mathrm{optimal}})$
13:    $\text{ari}\leftarrow\text{ComputeARI}(\text{feature\_clusters},\,\text{target\_clusters})$
14:    if $\text{ari}>\text{best\_ari}$ then
15:     $\text{best\_ari}\leftarrow\text{ari}$
16:     $\text{best\_features\_combination}\leftarrow\text{combination}$